• Suppose a firm has limited money expenditure of 800 birr to be spent on two factors labour and capital. The price of capital is per hours and the price of labour is birr 10 per hours.
  • A) draw the iso cost line
  • B) calculate the slope of the iso cost line
1
Expert's answer
2022-02-07T15:34:18-0500

Iso cost line

Qn. A

Qn. B

"slope=\frac{wage}{r}"

"Wage=10"

"r=20"

"slope=- \frac{10}{20}" "=-\frac{1}{2}"